OpenStudy (anonymous):

Which of the following equations represents an inverse variation? A. 6y = x B. 1/2y=1/x C. –x = y + 8 D. y = 1/4x

OpenStudy (anonymous):

award and medal

OpenStudy (anonymous):

didn't mean to close it

OpenStudy (welshfella):

with a direct variation as one variable increases or decreases the other variable increases or decreases eg y = 4 x y= 2, x = 1/2 y = 4 , x= 1

OpenStudy (anonymous):

c?

OpenStudy (welshfella):

for an inverse variation as y increases x decreases and vice versa

OpenStudy (anonymous):

@welshfella is it c

OpenStudy (welshfella):

yes c is one

OpenStudy (welshfella):

there is also another another inverse function is y = 1/x

OpenStudy (welshfella):

can you see a similar one to that ?

OpenStudy (anonymous):

yeah b right? which one would you choose if you had to? I feel like its c because it makes more sense

OpenStudy (welshfella):

c is correct but there is another one similar to the one i mentioned in my last post

OpenStudy (welshfella):

b is incorrect because if you invert both sides it gives x = 2y which is direct variation

OpenStudy (welshfella):

is d